# Adaptive Transfer Clustering: A Unified Framework

### v2

@@ D.3 Proof of Lemma 3
[…]
@@
[…]
@@ D.5 Proof of Lemma 7
$\geq\mathbb{P}_{1,1}\left(X_{0,i}<-K^{+}\right)\mathbb{P}_{1,1}\left(X_{1,i}>K^{+}\right)=\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\left(1-\Phi\left(\frac{-\mu+K^{+}}{\sigma}\right)\right).$
@@
$=\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{1,i}\geq K^{-}\right\}\right)$
@@
$=\frac{1}{\sqrt{2}}\frac{\sqrt{2/\pi}\exp\left(-\mu^{2}/\sigma^{2}\right)}{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}\cdot\frac{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}{\left(K^{-}+\mu\right)/\sigma+\sqrt{\left(K^{-}+\mu\right)^{2}/\sigma^{2}+8/\pi}}\Phi\left(-\sqrt{2}K^{-}/\sigma\right)$
@@
$\geq\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\Phi\left(\frac{\mu-K^{+}}{\sigma}\right)+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\left(1-2\Phi\left(-\frac{\sqrt{2}K^{-}}{\sigma}\right)\right)\vee\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}.$
@@ D.6 Proof of Lemma 8
Bound for $\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)$.
@@
Bound for $\mathbb{P}\left(\check{Z}_{0,i}^{\lambda}\neq\bar{Z}_{0,i}^{\lambda}\right)$.
@@ D.7 Proof of Lemma 9
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,\quad-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),\quad X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1-\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
@@
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,\quad-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),\quad X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
@@ D.8 Proof of Lemma 10
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$ $\leq\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}<\eta<-\sqrt{2}\frac{\mu}{\sigma}\left(1-\delta_{n}\right)\right)+\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}\left(1+\delta_{n}\right)<\eta<-\sqrt{2}\frac{\mu}{\sigma}\right)+n^{-C_{\sf par}}$
@@
$-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{1,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{2\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
@@
$V_{1,i}>\max\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma}Z^{*}_{0,i},$
@@
$V_{1,i}<\min\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}.$
@@ D.10 Proof of Lemma 12
$=\left(1+\vartheta\right)\vartheta_{1}\cdot\frac{1}{n}\sum_{i=1}^{n}X_{1,i}+\left(1+\frac{2}{\vartheta}\right)\left(1+\left(1+\vartheta\right)\vartheta_{1}\right)\frac{7\log\left(2/\delta\right)}{6n}+2\vartheta_{2}$ (by $\vartheta=\frac{2\upsilon}{1-\upsilon}$)
$\leq\frac{\left(1+\vartheta\right)\vartheta_{1}}{n}\sum_{i=1}^{n}X_{1,i}+\frac{7\log\left(2/\delta\right)}{2\vartheta n}\left(1+2\vartheta_{1}\right)+2\vartheta_{2}.$ (by $\vartheta<1$)

### v3

@@ D.3 Proof of Lemma 3
$\leq\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)+\gamma\left[\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}_{0,i}^{\infty}\right)+2\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\right].$
@@
$\leq\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)+\gamma\left[\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}_{0,i}^{\infty}\right)+2\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\right]$
@@ D.5 Proof of Lemma 7
$=\mathbb{P}_{1,1}\left(X_{0,i}<-\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},X_{1,i}>\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},{\cal E}_{\sf par}\right)$
$\geq\mathbb{P}_{1,1}\left(X_{0,i}<-K^{+}\right)\mathbb{P}_{1,1}\left(X_{1,i}>K^{+}\right)$
$=\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\left(1-\Phi\left(\frac{-\mu+K^{+}}{\sigma}\right)\right).$
@@
$=\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)$
$-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{1,i}\geq K^{-}\right\}\right).$
@@
$\mathbb{P}$ ($\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)$
@@
$=\frac{1}{\sqrt{2}}\frac{\sqrt{2/\pi}\exp\left(-\mu^{2}/\sigma^{2}\right)}{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}\cdot\frac{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}{\left(K^{-}+\mu\right)/\sigma+\sqrt{\left(K^{-}+\mu\right)^{2}/\sigma^{2}+8/\pi}}\Phi\left(-\frac{\sqrt{2}K^{-}}{\sigma}\right)$
@@
$\geq\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\Phi\left(\frac{\mu-K^{+}}{\sigma}\right)+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)$
$+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\left(1-2\Phi\left(-\frac{\sqrt{2}K^{-}}{\sigma}\right)\right)\vee\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}.$
@@ D.6 Proof of Lemma 8
Bound for $\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)$
@@
Bound for $\mathbb{P}\left(\check{Z}_{0,i}^{\lambda}\neq\bar{Z}_{0,i}^{\lambda}\right)$
@@ D.7 Proof of Lemma 9
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1-\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
@@
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
@@ D.8 Proof of Lemma 10
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$
$\leq\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}<\eta<-\sqrt{2}\frac{\mu}{\sigma}\left(1-\delta_{n}\right)\right)+\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}\left(1+\delta_{n}\right)<\eta<-\sqrt{2}\frac{\mu}{\sigma}\right)+n^{-C_{\sf par}}$
@@
$-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{1,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$-\frac{2\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
@@
$V_{1,i}>\max\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma}Z^{*}_{0,i},$
@@
$V_{1,i}<\min\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}.$
@@ D.10 Proof of Lemma 12
$=\left(1+\vartheta\right)\vartheta_{1}\cdot\frac{1}{n}\sum_{i=1}^{n}X_{1,i}+\left(1+\frac{2}{\vartheta}\right)\left(1+\left(1+\vartheta\right)\vartheta_{1}\right)\frac{7\log\left(2/\delta\right)}{6n}+2\vartheta_{2}$
$\leq\frac{\left(1+\vartheta\right)\vartheta_{1}}{n}\sum_{i=1}^{n}X_{1,i}+\frac{7\log\left(2/\delta\right)}{2\vartheta n}\left(1+2\vartheta_{1}\right)+2\vartheta_{2},$
where we’ve used $\vartheta=\frac{2\upsilon}{1-\upsilon}$ in the second equality and $\vartheta<1$ in the last inequality.